### v1

# Simplex volumes in hyperplane arrangements

Abstract
The Erdős’s distinct distance problem and unit distance problem are classical problems posed by Erdős [12] in 1946, which ask for the minimum number of distinct distances determined by $n$ points in $\mathbb{R}^{d}$ and the maximum number of pairs at a given distance that such a set can form, respectively. These problems have been studied for many decades, and a variety of interesting results have been obtained. Well-studied variants of these problems ask for the minimum number of distinct volumes of $d$-simplices determined by $d+1$ points among $n$ points in $\mathbb{R}^{d}$, and the maximum number of $d$-simplices with unit volume among $n$ points in $\mathbb{R}^{d}$. We propose the following dual setting of these problems and give partial results:
  • Determine the maximum number of unit-volume $d$-simplices determined by $n$ hyperplanes in $\mathbb{R}^{d}$.
  • Determine the maximum number of minimum/maximum-volume $d$-simplices determined by $n$ hyperplanes in $\mathbb{R}^{d}$.
  • Determine the maximum number $D_{d}(n)$ such that any arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$ in general position contains a subset of size $D_{d}(n)$ whose induced $d$-simplices have all distinct volumes.
Department of Applied Mathematics,
Faculty of Mathematics and Physics, Charles University, Czech Republic
koki@kam.mff.cuni.cz
1 Introduction
For $d\geq 2$, a set of points $X$ in $\mathbb{R}^{d}$ with $|X|\geq d+1$ is said to be in general position if no $d+1$ points from $X$ lie on the same (d1)(d-1)-dimensional hyperplane, and we say that an arrangement of hyperplanes $\mathcal{A}$ in $\mathbb{R}^{d}$ with $|\mathcal{A}|\geq d+1$ is in general position if every $d$ hyperplanes intersect at a single point and no $d+1$ have a common point.
Throughout this paper, we use $O_{d}$ to denote the standard big-$O$ notation where the constant may depend on $d$. We adopt similar conventions for $\Omega_{d}$ and $\Theta_{d}$.
1.1 Erdős’s distinct distance problem
In 1946, Erdős [12] posed the question: what is the minimum number $g_{d}(n)$ of distinct distances determined by $n$ points in $\mathbb{R}^{d}$. This question is known as Erdős’s distinct distance problem, and a variety of interesting results have been obtained. For the planar case $d=2$, Erdős showed an upper bound $g_{2}(n)=O(n/\sqrt{\log n})$ by considering the $\sqrt{n}\times\sqrt{n}$ square lattice. A lower bound for $g_{2}(n)$ has been improved many times [31, 3, 4, 40, 37, 43, 27]. The most recent bound is due to Guth and Katz [25], who showed that $g_{2}(n)=\Omega(n/\log n)$, which almost matches Erdős’s upper bound.
In higher dimensions, considering a $d$-dimensional grid $n^{1/d}\times\cdots\times n^{1/d}$, we obtain $g_{d}(n)=O_{d}(n^{2/d})$, and this bound is conjectured to be optimal. The best known lower bound to date is due to Solymosi and Vu [38] which is $g_{d}(n)=\Omega_{d}(n^{2/d-2/(d(d+2))})$.
A well-studied variant of this problem asks for the minimum number $g^{\prime}_{d}(n)$ of distinct $d$-dimensional volumes of $d$-simplices spanned by $n$ points in $\mathbb{R}^{d}$, assuming that not all $n$ points lie on a common hyperplane. One easily sees that $g^{\prime}_{d}(n)\leq\lfloor(n-1)/d\rfloor$ by taking $d$ sets of about $n/d$ equally spaced points on $d$ parallel lines passing through the vertices of a fixed (d1)(d-1)-simplex. Erdős, Purdy, and Straus [16] conjectured that this construction is tight for all sufficiently large $n$. Pinchasi [33] proved this conjecture for the case $d=2$, and Dumitrescu and Cs.Tóth [11] later showed that $g^{\prime}_{d}(n)=\Theta_{d}(n)$ for every $d$.
There is another related problem asking (A): what is the maximum number $h_{d}(n)$ such that any set of $n$ points in $\mathbb{R}^{d}$ contains a subset $S$ of size $h_{d}(n)$ whose induced $d$-simplices all have distinct $d$-volumes. The upper bound is again derived from the $n^{1/d}\times\cdots\times n^{1/d}$ grid, which gives $h_{2}(n)=O_{d}(n^{2/3})$ and $h_{d}(n)=O_{d}(n^{(d-1)/d})$ for $d\geq 3$. Conlon et al. [7] showed $h_{d}(n)=\Omega_{d}(n^{1/(2d+2)})$ for all $d\geq 2$.
1.2 Erdős’s unit distance problem
The unit distances problem asks: what is the maximum number $u_{d}(n)$ of unit distances that a set of $n$ points in $\mathbb{R}^{d}$ can form. For the case $d=2$, Erdős [12] obtained the bounds $n^{1+c/\log\log n}<u_{2}(n)=O(n^{3/2})$ and conjectured that his lower bound is tight. The problem remains wide open; the best known upper bound $u_{2}(n)=O(n^{4/3})$ is due to Spencer, Szemerédi, and Trotter [39]. Despite numerous efforts [5, 40, 32], no further improvements have been found. In $3$-dimensions, the best known bounds are $u_{3}(n)=O(n^{3/2})$ [26, 44] and $u_{3}(n)=\Omega(n^{4/3}\log\log n)$ [13]. For $d\geq 4$, we have $u_{d}(n)=\Theta_{d}(n^{2})$ (see [30, 14]).
This problem also has a well-studied direction of generalization. In 1967, Oppenheim asked (B): what is the maximum number of unit area triangles determined by $n$ points in the plane? A simple lattice construction $\sqrt{n}\times\sqrt{n}$ again gives a lower bound $\Omega(n^{2}\log\log n)$ [15]. Currently, the best known upper bound is $O(n^{2+2/9})$, due to Raz and Sharir [34]. Dumitrescu and Cs.Tóth [11] extended this problem to three dimensions and obtained the following result: the number of unit volume tetrahedra ($3$-simplices) determined by $n$ points in $\mathbb{R}^{3}$ is $O(n^{7/2})$, and there is an $n$-point set achieving $\Omega(n^{3}\log\log n)$.
Related problems (C): are to determine the number of triangles of minimum/maximum area determined by a set of $n$ points studied by Braß, Rote, and Swanepoel [2]. They proved that the maximum number of triangles of maximum area determined by $n$ points in the plane is exactly $n$, and the maximum number of triangles of minimum area determined by $n$ points in the plane is $\Theta(n^{2})$.
Dumitrescu and Cs.Tóth [11] further generalized this problem and obtained the following results: The number of tetrahedra of minimum (non-zero) volume spanned by $n$ points in $\mathbb{R}^{3}$ is at most $2n^{3}/3-O(n^{2})$, and there is a configulation of $n$-points achieving $3n^{3}/16-O(n^{2})$. Furthermore, they proved that the maximum number of $d$-simplices of minimum volume spanned by $n$ points in $\mathbb{R}^{d}$ is $O_{d}(n^{d})$ for all $d$.
1.3 Dual setting
In this paper, we investigate the following dual setting of (A), (B) and (C).
  1. 1.
    The following is a dual setting of (A):
    Question 1.
    What is the maximum number $D_{d}(n)$ such that any arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$ in general position contains a subset of size $D_{d}(n)$ whose induced $d$-simplices have all distinct volumes?
    For the case $d=2$, Damásdi et al. [9] obtained a lower bound of $n^{1/5}$ under the stronger assumption that there are no six lines that are tangent to a common conic (note that five lines always have a common tangent conic).
  2. 2.
    The problem of generalizing (B) to higher dimensions is to determine the maximum number of unit volume $d$-simplices formed by n points in $\mathbb{R}^{d}$. For this problem, we consider the following dual version.
    Question 2.
    What is the maximum number $f_{d}(n)$ of unit-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$?
    For the case $d=2$, Damásdi et al. [9] proved that $f_{2}(n)=O(n^{9/4+\varepsilon})$ for every fixed $\varepsilon>0$, and $f_{2}(n)=\Omega(n^{2})$. Their proof uses the incidence theorem between points and algebraic curves, proved by Sharir and Zahl [36].
  3. 3.
    We propose the following variants for the higher-dimensional generalization of (B) (that is, to determinie the number of $d$-simplices of minimum/maximum volume determined by a set of $n$ points in $\mathbb{R}^{d}$).
    Question 3.
    What is the maximum number $m_{d}(n)$ of minimum-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$ ?
    For the case $d=2$, Damásdi et al. [9] proved that $m_{2}(n)=\Theta(n^{2})$.
    Question 4.
    What is the maximum number $M_{d}(n)$ of maximum-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$ ?
    As we mentioned above, with respect to the occurrence of the maximum area determined by $n$ points in the plane, it is known that the maximum number is exactly $n$. However, Damásdi et al. [9] showed that this phenomenon does not occur in the dual setting: $M_{2}(n)>7n/5-O(1)$. They also showed an upper bound $M_{2}(n)\leq n(n-1)/3$.
1.4 Our results
We provide partial answers to Questions 1-4, especially for $d\geq 3$.
Theorem 1.1.
For the maximum number $f_{d}(n)$ of unit-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
$f_{d}(n)=O_{d}\bigl(n^{\,d+1-\tfrac{d}{d^{d}+1}}\bigr).$
Theorem 1.2.
$f_{d}(n)=\Omega_{d}(n^{d}).$
Theorem 1.3.
For the maximum number $m_{d}(n)$ of minimum-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
$m_{d}(n)=\Theta_{d}(n^{d}).$
Theorem 1.4.
For the maximum number $M_{3}(n)$ of maximum-volume tetrahedra formed by $n$ planes in $\mathbb{R}^{3}$, we have
$M_{3}(n)>\frac{7}{6}n-O(1).$
Theorem 1.5.
$D_{2}(n)<<n(\log n)^{-c}$ for some absolute constant $c>0$, and for $d\geq 3$, $D_{d}(n)<<ne^{-(\log\log n)^{c_{d}}}$ for some absolute constant $c_{d}\in(0,1)$.
Theorem 1.5 gives the a partial answer for Problem 5.4 and Problem 5.5 in [9].
The rest of the paper is organized as follows. We will prove Theorem 1.1, 1.2 in the next section. In Section 3 we will prove Theorem 1.3. Section 4 is devoted to proving Theorem 1.4. Finally, in Section 5, we will give a proof of Theorem 1.5.
2 Unit volume $d$-simplices
In this section, we consider the number of $d$-simplices with unit volume determined by an arrangement of $n$ hyperplanes, namely to investigate the magnitude of $f_{d}(n)$. We have a trivial upper bound $O_{d}(n^{d+1})$ by selecting generic $d+1$ hyperplanes from the arrangement. The following theorem provides a slightly better upper bound for $f_{d}(n)$. See 1.1 First, we see the following proposition.
Proposition 2.1.
Let $H_{1},\ldots,H_{d}$ be hyperplanes in $\mathbb{R}^{d}$ whose intersection is a single point. For a fixed coordinate system $(x_{1},\ldots,x_{d})$, there exist exactly two hypersurfaces of degree $d$ obtained by an (invertible) affine transformation of the hypersurface $\prod_{i=1}^{d}x_{i}=c$ for some $c\neq 0$ such that each hyperplane tangent to them and $H_{1},\ldots,H_{d}$ form a $d$-simplex with unit volume.
Proof.
Note that an (invertible) affine transformations preserve hyperplanes, hypersurfaces of degree $d$ and ratios of $d$-volumes. Therefore, by applying an appropriate affine transformation, we may set $H_{i}=\{x_{i}=0\}$ for $i\in\{1,\ldots,d\}$. Now we fix $c>0$ and consider the hypersurface $\prod_{i=1}^{d}x_{i}=c$. Then the equation of the hyperplane $h_{p}$ tangent to this surface at a point $p=(p_{1},\ldots,p_{d})$ is given by $\sum_{i=1}^{d}x_{i}/p_{i}=d$. The hyperplane $h_{p}$ and $H_{1},\ldots,H_{d}$ form a $d$-simplex $\Delta_{p}$ with vertices $(0,\ldots,0,dp_{i},0,\ldots,0)$ $(1\leq i\leq d)$, where $dp_{i}$ is in the $i$-th coordinate. Hence, the volume of $\Delta_{p}$ is
$\frac{1}{d!}\cdot\prod_{i=1}^{d}dp_{i}=\frac{d^{d}}{d!}|c|,$
which is a constant depending only on the dimension $d$. Similarly for $c<0$, any tangent hyperplane of $x_{1}x_{2}\cdots x_{d}=c$ and $H_{1},\ldots,H_{d}$ form a $d$-simplex with the same volume $d^{d}|c|/d!$. ∎
Proof of Theorem 1.1.
The proof is based on [10]. Let $\mathcal{H}$ be an arrangement of $n$ hyperplanes. Each of the two hypersurfaces given in the above proposition has $2^{d-1}$ connected branches ($2^{d}$ connected branches in total). Let $\sum\nolimits_{d}$ be a set of all branches obtained by selecting $d$ hyperplanes intersecting at a single point from $\mathcal{H}$. Then, we have $|\sum\nolimits_{d}|\leq 2^{d}\binom{n}{d}$. Note that for each branch $\sigma\in\sum\nolimits_{d}$, the dual of the set of all hyperplanes tangent to $\sigma$ is a subset of some hypersurface of degree $d$, and the number of its connected component is: two when there exists a hyperplane tangent to $\sigma$ parallel to the last axis, and one otherwise. By dual transformation in $\mathbb{R}^{d}$, we mean the bijection between points $p=(p_{1},p_{2},\ldots,p_{d-1},p_{d})\in\mathbb{R}^{d}$ and hyperplanes $p^{*}:x_{d}=p_{1}x_{1}+p_{2}x_{2}+\cdots+p_{d-1}x_{d-1}+p_{d}$ in $\mathbb{R}^{d}$. Now we consider two bipartite graphs $G_{1},G_{2}$ with vertex set $\mathcal{H}\times\sum\nolimits_{d}$ as follows. If $H\in\mathcal{H}$ is tangent to $\sigma\in\sum\nolimits_{d}$ from above (in the sense of the last axis), we define an edge $(H,\sigma)\in\mathcal{H}\times\sum\nolimits_{d}$ in $G_{1}$. Similarly, If $H\in\mathcal{H}$ is tangent to $S\in\sum\nolimits_{d}$ from below, we define an edge $(H,\sigma)\in\mathcal{H}\times\sum\nolimits_{d}$ in $G_{2}$. The edges of $G_{1}$ and $G_{2}$ correspond to all hyperplane-hypersurface tangencies. Thus, to obtain an upper bound for $f_{d}(n)$, it suffices to bound the number of edges $E(G_{1})$ in $G_{1}$ and the number of edges $E(G_{2})$ in $G_{2}$. Observe that neither $G_{1}$ nor $G_{2}$ contains $K_{d^{d}+1,d}$. Indeed, if $G_{1}$ contains $K_{d^{d}+1,d}$, then there would be $d^{d}+1$ hyperplanes in $\mathcal{H}$ tangent to $d$ branches from above. Considering this situation in a dual space, meaning that $d$ branches intersect at $d^{d}+1$ points, which contradicts the higher-dimensional Bézout’s theorem (refer to [42]). The same argument applies to $G_{2}$. Therefore, by the Kövári–Sós–Turán theorem,
$|E(G_{1})|,|E(G_{2})|=O\bigl(n\cdot n^{d(1-\frac{1}{d^{d}+1})}+n^{d}\bigr)=O\bigl(n^{\,d+1-\tfrac{d}{d^{d}+1}}\bigr).$
Clearly, a lower bound for $m_{d}(n)$ gives a lower bound for $f_{d}(n)$. See the result in the next section.
3 Minimum volume $d$-simplices
In this section, we prove Theorem 1.3. Let us recall the statement. See 1.3
Kobon Fujimura asked in his book (called ”The Tokyo puzzle” [18]) that what is the largest number $K(n)$ of non-overlapping triangles whose sides lie on an arrangement of $n$ lines (Grünbaum [24] had studied a similar problem slightly earlier than Fujimura, but only in the case of arrangements in the projective plane). A first upper bound was established by Saburo Tamura, which is $K(n)\leq\lfloor\frac{n(n-2)}{3}\rfloor$. In 2007, Clément and Bader [6] found a slightly improved upper bound:
$K(n)\leq\begin{cases*}\lfloor\frac{n(n-2)}{3}\rfloor-1&if $n\equiv 0,2\pmod{6}$,\\ \lfloor\frac{n(n-2)}{3}\rfloor&otherwise.\end{cases*}$
For even $n$, Bartholdi et al. [1] provided a slightly sharper upper bound of $\lfloor\frac{n}{3}(n-\frac{7}{3})\rfloor$. On the other hand, Füredi and Palásti [19] having constructed an arrangement of $n$ lines which contains $\lfloor\frac{n(n-3)}{3}\rfloor$ non-overlapping triangles, which gives $K(n)\geq\lfloor\frac{n(n-3)}{3}\rfloor$.
We now consider a higher-dimensional analogue. Let $K_{d}(n)$ be the largest number of nonoverlapping $d$-dimensional simplices whose facets lie on an arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$. Note that $K(n)=K_{2}(n)$.
Proposition 3.1.
For any $d\geq 3$, and $n\geq d+1$,
$K_{d}(n)\leq\frac{2}{(d+1)!}n^{d}+O_{d}(n^{d-1}).$
The proof is presented in the appendix. For $d=2$, it immediately follows that $m_{2}(n)\leq K_{2}(n)$ from the fact that cutting a triangle with a single line always creates a smaller triangle in its interior ([9]). However, unfortunately, for $d\geq 3$, the same argument does not hold. Indeed, a $d$-simplex cut by a hyperplane creates a new $d$-simplex with volume smaller than the original in the interior only if the hyperplane separates one vertex from all the others. See Figure 1 for the case $d=3$. We provide an upper bound for $m_{d}(n)$ by a different approach.
Figure 1: Left: A cut that creates a $3$-simplex in the interior of a 33-simplex. Right: A cut that does not create a $3$-simplex in the interior.
Proof of Theorem 1.3.
Upper bound: We begin by demonstrating the three-dimensional case. Let $\mathcal{H}$ be an arrangement of $n$ planes in $\mathbb{R}^{3}$, and let $V_{0}$ be the volume of a tetrahedron with minimum volume formed by $\mathcal{H}$. We denote by $\tau(H_{1},H_{2},H_{3},H_{4})$ the tetrahedron determined by $H_{1},H_{2},H_{3},H_{4}$, and let $\tilde{H_{i}}\subset H_{i}\,\,(i=1,2,3,4)$ be its faces.
Consider $H_{1},H_{2},H_{3}\in\mathcal{H}$ intersecting at a single point. These divide $\mathbb{R}^{3}$ into $8$ connected components. Let $R$ be one of them. Then by Proposition 2.1, there exists a connected branch $\sigma_{R}$ of a surface in the region $R$ such that for any point $p$ on $\sigma_{R}$, the tangent plane $T_{p}(\sigma_{R})$ satisfies that the tetrahedron $\tau(H_{1},H_{2},H_{3},T_{p}(\sigma_{R}))$ has volume $V_{0}$. Moreover, by fixing a coordinate system $(x,y,z)$ and applying an appropriate affine transformation, we may assume $H_{1}=\{z=0\}$, $H_{2}=\{y=0\}$, $H_{3}=\{x=0\}$. Then the connected branch corresponding to $\sigma_{R}$ is $C_{R}:xyz=2V_{0}/9\,\,(x,y,z>0)$. For a point $p$ on $C_{R}$, let $T_{p}(C_{R})$ denote its tangent plane.
Claim 1.
$|\{p\in C_{R}:T_{p}(C_{R})\in\mathcal{H}\}|\leq 1$, that is, there is at most one $H\in\mathcal{H}$ such that the tetrahedron $\tau(H_{1},H_{2},H_{3},H)$ has a minimum volume in $R$.
Proof.
[…]
Figure 2:
Suppose there are two planes $T,T^{\prime}\in\mathcal{H}$ such that for some $p,p^{\prime}\in C_{R}$, we have $T=T_{p}(C_{R}),\,\,T^{\prime}=T_{p^{\prime}}(C_{R})$. Then the interior of the face $\tilde{T}$ of the tetrahedron $\tau=\tau(H_{1},H_{2},H_{3},T)$ intersects $T^{\prime}$. Without loss of generality, we may assume that $T^{\prime}$ does not intersect only $H_{3}\cap\tilde{T}$ among the three edges of $\tilde{T}$. Let $s_{i}$ denote the intersection point of $T^{\prime}$ with $H_{i}\cap T$ for $i\in\{1,2\}$. The vertex $H_{1}\cap H_{2}\cap T$ of $\tau$ can be represented as $S=(s,0,0)$ for some $s>0$. Also, the vertex $H_{1}\cap H_{2}\cap H_{3}$ of $\tau$ is the origin $O$. From the positions of $s_{1}$ and $s_{2}$, it follows that $T^{\prime}$ must intersect the open half-line $\{(t,0,0):t>0\}(\subset H_{1}\cap H_{2})$. Let $H_{1}\cap H_{2}\cap T^{\prime}=\{(u,0,0)\}$. We may assume $u>s$. Indeed, otherwise the tetrahedron $Sus_{1}s_{2}$ would be contained in the interior of $\tau$, contradicting the minimality of the volume of $\tau$. When $u>s$, $T^{\prime}$ intersects $\tilde{H_{1}}\cap\tilde{H_{3}}$ at some point $t_{1}$ and $\tilde{H_{2}}\cap\tilde{H_{3}}$ at some point $t_{2}$. Then the tetrahedron $Out_{1}t_{2}$ contains the tetrahedron $Sus_{1}s_{2}$ in its interior. However, the tetrahedron $Out_{1}t_{2}$ is $\tau(H_{1},H_{2},H_{3},T^{\prime})$, contradicting the minimality of volume. See Figure 2. ∎
By the same argument for the other connected components besides $R$, we find that there are at most $8$ planes $H\in\mathcal{H}$ such that $\tau(H_{1},H_{2},H_{3},H)$ has minimum volume $V_{0}$. Therefore, the number of tetrahedra with minimum volume in $\mathcal{H}$ is at most $8\binom{n}{3}$.
It turns out that an even simpler approach works for higher dimensions as follows. As in the three-dimensional case, we consider hyperplanes $H_{1}=\{x_{1}=0\},\ldots,H_{d}=\{x_{d}=0\}$ in $\mathbb{R}^{d}$ and a connected branch $C_{R}$ of a hypersurface of degree $d$ belonging to the region $R=\mathbb{R}_{\geq 0}^{d}$ which satisfy the property of Proposition 2.1 with minimum volume. We fix two tangent hyperplanes $T,T^{\prime}$ of $C_{R}$, and denote the $d$-simplices they form by $\tau=\tau(H_{1},\ldots,H_{d},T)$ and $\tau^{\prime}=\tau(H_{1},\ldots,H_{d},T^{\prime})$, respectively. Now, by a well-known result of Shannon [35], the arrangement of $d+2$ hyperplanes $\{H_{1},\ldots,H_{d},T,T^{\prime}\}$ has two $d$-simplicial cells $\sigma_{1},\sigma_{2}$. Clearly, $T^{\prime}$ must passes through the interior of $\tau$ and divides $\tau$ into two regions $R_{1},R_{2}$. We may assume that neither $R_{1}$ nor $R_{2}$ is a $d$-simplicial cell (otherwise this would contradict the minimality of the volume of $\tau$). Therefore, neither $\sigma_{1}$ nor $\sigma_{2}$ intersects $\tau$. Then, $\tau^{\prime}$ can be expressed as the union of one of $R_{1},R_{2}$ and one of $\sigma_{1},\sigma_{2}$. This contradicts the minimality of the volume of $\tau^{\prime}$. Hence, the number of minimum $d$-volume $d$-simplices is at most $2^{d}\binom{n}{d}$.
Lower bound: Let us denote by $S_{d}$ the set of all permutations of $\{1,\ldots,d\}$ and let $\pi\in S_{d}$ be a permutation. The Coxeter–Freudenthal–Kuhn (CFK) triangulation of the unit hypercube $0,1]^{d}$ is a triangulation of $0,1]^{d}$ into $d$-simplices $\sigma_{\pi}=[v_{0}(\pi),\ldots,v_{d}(\pi)]$, whose vertices are given by
$v_{0}(\pi)=(0,\ldots,0),\quad v_{i}(\pi)=v_{i-1}(\pi)+e_{\pi(i)}\quad\text{for }i=1,\ldots,d,$
where $e_{1},\ldots,e_{d}$ is a standard basis of $\mathbb{R}^{d}$. Note that the points $v_{0}(\pi),\ldots,v_{d}(\pi)\in\mathbb{R}^{d}$ are vertices of $0,1]^{d}$ and the CFK triangulation decomposes $0,1]^{d}$ into $d!$ $d$-simplices with volume $1/d!$. It is easy to observe that this decomposition is realized by $\binom{d}{2}$ hyperplanes
$\{x_{p}-x_{q}=0:1\leq p<q\leq d\}.$
See Figure 3 and refer to [8], [17], [28] for further details
(1,1,1)(1,1,1)(0,0,0)(0,0,0)(0,0)(0,0)$(1,1)$
Figure 3: The CFK triangulation of $0,1]^{2}$ and $0,1]^{3}$.
Now, we consider $d(n+1)$ hyperplanes given by
$x_{i}=k\quad(i\in\{1,\ldots,d\},\,\,k\in\{0,\ldots,n\}),$
which form an arrangement that decomposes the hypercube $0,n]^{d}$ into $n^{d}$ unit hypercubes. Furthermore, by adding $\binom{d}{2}(2n-1)$ hyperplanes
$\{x_{p}-x_{q}=t:1\leq p<q\leq d,\,t\in\mathbb{Z},\,|t|\leq n-1\},$
We can see that each unit hypercube admits the CFK triangulation (this is called K1 CFK triangulation). The size of the resulting hyperplane arrangement is
$d(n+1)+\binom{d}{2}(2n-1)=d^{2}n+\frac{d(3-d)}{2},$
and the number of (minimum volume) $d$-simplices is
$d!\cdot n^{d},$
which gives us a lower bound $\Omega_{d}(n^{d})$.
4 Maximum volume tetrahedra
In this section, we show a lower bound for $M_{3}(n)$. A linear lower bound for $M_{3}(n)$ can be easily obtained using $3$ planes together with n3n-3 tangent planes to a hypersurface having them as asymptotic planes. By generalizing the argument in [9] to 3 dimensions, we obtain the following lower bound. See 1.4 The proposition they proved for their construction giving $M_{2}(n)>6n/5-O(1)$ can be extended as follows.
Proposition 4.1.
If there are no parallel planes in the arrangement, then we can find a rectangular box $ABCDEFGH$ such that adding any plane to the arrangement that intersects the segments $AB$, CDCD, EFEF and $GH$ creates no new maximum volume tetrahedra.
For a plane arrangement $\mathcal{A}$, let $T(\mathcal{A})$ denote the number of maximum volume tetrahedra in $\mathcal{A}$.
$A_{1}$$B_{1}$$D_{1}$$C_{1}$$H_{1}$$G_{1}$$E_{1}$$F_{1}$$A_{2}$$B_{2}$$C_{2}$$D_{2}$$E_{2}$$F_{2}$$G_{2}$$H_{2}$
Figure 4: Rectangular boxes
Figure 5: A star-shaped badge
Proposition 4.2 (Generalization of Proposition 4.2 in [9]).
If $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are arrangements of planes which contains no parallel planes, then there exists affine transformations $\varphi,\psi$ such that
$T(\varphi(\mathcal{H}_{1})\cup\psi(\mathcal{H}_{2}))\geq T(\mathcal{H}_{1})\cup T(\mathcal{H}_{2})+2.$
Proof.
In an arrangement $\mathcal{H}_{1}$, there exists a rectangular box satisfying Proposition 4.1 with vertices $A_{1},B_{1},C_{1},D_{1},E_{1},F_{1},G_{1},H_{1}$, and similarly in $\mathcal{H}_{2}$, there exists one with vertices $A_{2},B_{2},C_{2},D_{2},E_{2},F_{2},G_{2},H_{2}$. By applying appropriate affine transformations to each, we can arrange $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ so that the volumes of their respective maximum-volume tetrahedra are equal and the two rectangular boxes cross each other as shown in the Figure 5. Let $\varphi$ denote the affine transformation applied to $\mathcal{H}_{1}$ and $\psi$ the affine transformation applied to $\mathcal{H}_{2}$.
As the next step, we translate the planes of $\psi(\mathcal{H}_{2})$ to create a new maximum-volume tetrahedron $S$. If exactly one of the four planes defining $S$ comes from a plane of $\varphi(\mathcal{H}_{1})$, let this plane be $H$. By moving the entire $\psi(\mathcal{H}_{2})$ along $H$, we can create another new maximum-volume tetrahedron. Similarly, if exactly three planes come from planes of $\varphi(\mathcal{H}_{1})$, we can move $\varphi(\mathcal{H}_{1})$ along one plane of $\psi(\mathcal{H}_{2})$. Finally, if exactly two planes come from planes of $\varphi(\mathcal{H}_{1})$, by moving $\psi(\mathcal{H}_{2})$ along the line of intersection of those two planes, we can create a new maximum-volume tetrahedron. ∎
We observe that the plane arrangement $\mathcal{S}$ forming a star-shaped badge as shown in the Figure 5 contains $6$ planes and $5$ tetrahedra of maximum volume. Starting from $\mathcal{S}$, we repeatedly add copies of $\mathcal{S}$ so as to satisfy the above property. This allows us to construct a planar arrangement of $6k$ planes containing 7k27k-2 tetrahedra of maximum volume for $k\geq 1$.
We are not aware of any construction providing a non-trivial lower bound for $M_{d}(n)$ for $d\geq 4$.
5 Hyperplanes defining distinct volume $d$-simplices
In this section, we show that for all $d\geq 2$, $D_{d}(n)$ grows strictly slower than $n$. See 1.5
Let $n]=\{1,\ldots,n\}$ and denote $r_{k}(n)$ to be the size of the largest $S\subseteq[n]$ such that $S$ contains no $k$ elements in arithmetic progression ($k$-AP). Here is the key proposition.
Proposition 5.1.
$D_{d}(n)<r_{d+2}(n)$ for all $d\geq 2$.
Determining the value of $r_{k}(n)$ is one of the fundamental problems in arithmetic combinatorics, and both upper and lower bounds have been progressively improved by many researchers. In 2009, Green and Tao [22, 23] proved that there exists a constant $c>0$ such that $r_{4}(n)\ll n(\log n)^{-c}$. For $k\geq 4$, Leng et al. [29] recently obtained an upper bound $r_{k}(n)<<ne^{-(\log\log n)^{c_{k}}}$ for some $c_{k}\in(0,1)$. See [41, 20, 21] for further details. Theorem 1.5 can be immediately derived from these bounds and Proposition 5.1. Therefore, any improvement of upper bound for $r_{k}(n)$ automatically improves that for $D_{d}(n)$.
Proof of Proposition 5.1.
We first consider the case $d=2$. Define $\mathcal{L}=\{l_{1},\dots,l_{n}\}$ as the set of lines corresponding to $n$ consecutive edges of a regular $N$-gon $(N>n)$, arranged in clockwise order. We choose $N$ sufficiently large so that any four lines in $\mathcal{L}$ form two triangles. By the definition of $r_{k}(n)$, every subset of $\mathcal{L}$ of size $r_{4}(n)+1$ always contains four lines $\{l_{i_{1}},l_{i_{2}},l_{i_{3}},l_{i_{4}}\}$ such that $\{i_{1},i_{2},i_{3},i_{4}\}\subseteq[n]$ is a $4$-terms arithmetic progression. One can easily see that these four lines then form two congruent triangles.
We generalize this construction to higher dimensions. Assume $d$ is even. Write $\mathbb{R}^{d}$ as the orthogonal sum of a two-dimensional subspace, thus
$\mathbb{R}^{d}=V_{1}\oplus\cdots\oplus V_{d/2}.$
For each $j$, let $C_{j}$ denote the unit circle centered at the origin in $V_{j}$. We choose angle parameters $\theta_{1},\ldots,\theta_{d/2}>0$ so that for each $j$, one has $n\theta_{j}\ll\pi/2,\,\text{and}\,\,\theta_{j}\neq\theta_{j^{\prime}}\ \text{for }j\neq j^{\prime}.$ For each $i\in\{0,\ldots,n\}$ and each $j\in\{1,\ldots,d/2\}$, we define $P_{i,j}=(\cos(i\theta_{j}),\sin(i\theta_{j}))\in V_{j}$ (viewed in the standard coordinates of $\mathbb{R}^{d}$, the point $P_{i,j}$ has nonzero coordinates only in its $V_{j}$-component, which is $(\cos(i\theta_{j}),\sin(i\theta_{j}))$, while all other components are zero). For each $i\in\{1,\ldots,n\}$, let $H_{i}$ denote the affine hull of
$\bigcup_{1\leq j\leq d/2}\{P_{i-1,j},P_{i,j}\}.$
This has dimension d1d-1. Let $\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ denote the resulting hyperplane arrangement. One can verify that $\mathcal{H}$ is in general position for almost all choices of the angle parameters $\theta_{1},\ldots,\theta_{d/2}$. For any positive integer $D$, we define the map $T_{D}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ as
$T_{D}|_{V_{j}}=R_{D\theta_{j}}\,(j=1,\ldots,d/2),$
where $R_{D\theta_{j}}$ is the rotation by angle $D\theta_{j}$ on $V_{j}$. One easily sees that $T_{D}(P_{i,j})=P_{i+D,j}$ (for $i+D\leq n$), and therefore $T_{D}(H_{i})=H_{i+D}.$ Since the Jacobian matrix of $T_{D}$ is
$J_{T_{D}}=\begin{pmatrix}R_{D\theta_{1}}&0&\cdots&0\\ 0&R_{D\theta_{2}}&\ddots&\vdots\\[2.0pt] \vdots&\ddots&\ddots&0\\[2.0pt] 0&\cdots&0&R_{D\theta_{d/2}}\end{pmatrix}$,
$\det J_{T_{D}}=\prod_{j=1}^{d/2}\det R_{D\theta_{j}}=1.$ Thus, the map $T_{d}$ is an affine volume-preserving map.
Now let $i_{1}<i_{2}<\cdots<i_{d+2}$ be a $(d+2)$-arithmetic progression in $[n]$ with common difference $L$. As the arrangement $\mathcal{H}$ is in general position, the hyperplanes $\{H_{i_{1}},\ldots,H_{i_{d+1}}\}$ form a unique $d$-simplex $S$, and the hyperplanes $\{H_{i_{2}},\ldots,H_{i_{d+2}}\}$ form a unique $d$-simplex $S^{\prime}$. Since $T_{L}(H_{i_{k}})=H_{i_{k+1}}$, we have $T_{L}(S)=S^{\prime}$ which means that $S$ and $S^{\prime}$ have the same volume. Therefore, by the definition of $r_{k}(n)$, every subfamily of $\mathcal{H}$ of size (at least) $r_{d+2}(n)+1$ always contain two $d$-simplices with the same volume.
Now, we assume $d$ is odd, we decompose $\mathbb{R}^{d}$ into an orthogonal direct sum of (d1)/2(d-1)/2 two-dimensional vector spaces and one one-dimensional vector space as follows:
$\mathbb{R}^{d}=V_{1}\oplus\cdots\oplus V_{(d-1)/2}\oplus L.$
First, we place $n$ equally spaced points $Q_{i}=(0,\ldots,0,i)$ ($i\in\{1,\ldots,n\}$) on $L$. Choose angular parameters $\theta_{1},\ldots,\theta_{(d-1)/2}>0$ such that $\theta_{j}\neq\theta_{j^{\prime}}$ for $j\neq j^{\prime}$. For each $j\in\{1,\ldots,(d-1)/2\}$, consider the helical curve wrapping around $L$ in the $3$-dimensional vector space $V_{j}\oplus L$:
$\gamma_{\theta_{j}}(t)=(0,\ldots,0,\underbrace{\cos(t\theta_{j})}_{2j-1},\underbrace{\sin(t\theta_{j})}_{2j},0,\ldots,0,t)$
and we place $(n+1)$ equally spaced points on $\gamma_{\theta_{j}}(t)$:
$P_{i,j}=(0,\ldots,0,\underbrace{\cos(i\theta_{j})}_{2j-1},\underbrace{\sin(i\theta_{j})}_{2j},0,\ldots 0,i),\quad i\in\{0,\ldots,n\}.$
Define $H_{i}$ as the affine hull of
$\bigcup_{1\leq j\leq(d-1)/2}\{P_{i-1,j},P_{i,j}\}\cup\{Q_{i}\},$
which has dimension d1d-1. By taking $\theta_{1},\ldots,\theta_{(d-1)/2}$ appropriately, we obtain the hyperplane arrangement $\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ which is in general position. For any positive integer $D$, one can define an affine volume-preserving map $T_{D}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ by
$T_{D}|_{V_{j}}=R_{D\theta_{j}},\,T_{D}|_{L}:m\mapsto m+D.$
Observe that $T_{D}(P_{i,j})=P_{i+D,j},\,T_{D}((0,\ldots,0,i))=(0,\ldots,0,i+D)$ (for $i+D\leq n$), hence $T_{D}(H_{i})=H_{i+D}.$ The same argument as in the previous case shows that any subfamily of $\mathcal{H}$ with size $r_{d+2}(n)+1$ contains two $d$-simplices with the same volume.
Remark 1.
As previouly mentioned, Damásdi et al. [9] obtained the lower bound with some additional assumptions: if there are no six lines that are tangent to a common conic (observe that $5$ lines always have a common tangent conic), then an arrangement of $n$ lines always contains a subset of size $n^{1/5}$ whose induced triangles have all distinct volumes. Their argument can be applied to general dimensions, and the following can be shown. If there are no $d(d+3)/2$ hyperplanes that are tangent to a common hypersurface, then then an arrangement of $n$ hyperplanes always contains a subset of size $n^{\frac{1}{2d+1}}$ whose induced $d$-simplices have all distinct volumes. On the other hand, We are not aware of any lower bounds for $D_{d}(n)$ for $d\geq 2$.
Acknowledgement. I would like to thank Pavel Valtr for helpful suggestions. The work was supported by grant no. 23-04949X of the Czech Science Foundation (GAČR).
Appendix
Proof of Proposition 3.1.
It suffices to show that
$K_{d}(n)\leq\frac{n}{d+1}K_{d-1}(n-1)$
holds for all $d\geq 3$. Suppose there are two parallel hyperplanes in an arrangement. If we perturb one of them by slightly rotating its normal vector, the number of $d$-simplicial cells in the arrangement does not decrease. Hence, we may assume that no two hyperplanes are parallel.
Consider an arrangement $\mathcal{H}$ of $n$ hyperplanes, and fix one hyperplane $H$ of $\mathcal{H}$. Then, restricting the remaining to $H$ yields an arrangement $\mathcal{S}$ of (d2)(d-2)-dimensional hyperplane in $H(\cong\mathbb{R}^{d-1})$. Since no two in $\mathcal{H}$ are parallel, no two in $\mathcal{S}$ are parallel either. The maximum number of (d1)(d-1)-simplicial cells in $\mathcal{L}$ is $K_{d-1}(n-1)$. Therefore, the total number of (d1)(d-1)-simplicial cells in $\mathcal{A}$ is at most $nK_{d-1}(n-1)$.
We now divide into two cases. First assume that no $d+1$ hyperplanes intersect at a single point. In this case, No two $d$-simplicial cells share a common (d1)(d-1)-simplicial face. Otherwise, at least one of the $d+1$ vertices of the common face intersects at least $d+1$ hyperplanes. See Figure 7. Therefore,
$(d+1)K_{d}(n)\leq nK_{d-1}(n-1).$
Next, assume that $d+1$ (or more) hyperplanes intersect at a single point. If a $(d+1)$-th hyperplane is added through an existing intersection point, the number of (d1)(d-1)-simplicial faces decreases by $d+1$. On the other hand, two $d$-simplicial cells sharing a common (d1)(d-1)-simplicial face can be created only by adding a new hyperplane passing through an existing point. See Figure 7. Therefore,
$(d+1)K_{d}(n)<nK_{d-1}(n-1).$
Figure 6:
Figure 7: In the case $d=3$: If four planes intersect at a single point, adding a fourth plane through an existing point decreases the number of triangles by four, with at most two of them are saved.