### v2

# Simplex volumes in hyperplane arrangements

Abstract
The distinct distances problem and the unit distance problem, posed by Erdős [erdos1946] in 1946, are classical problems in combinatorial geometry. They ask for the minimum number of distinct distances determined by $n$ points in $\mathbb{R}^{d}$ and the maximum number of pairs at unit-distance that such a set can determine, respectively. These problems have been studied extensively for several decades, leading to numerous deep results. We study the following dual setting:
  • Determine the maximum number of unit $d$-dimensional $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$.
  • Determine the maximum number of minimum (or maximum) $d$-dimensional $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$.
  • Determine the maximum number $D_{d}(n)$ such that any arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$ in general position contains a subset of size $D_{d}(n)$ for which all induced $d$-dimensional simplices have distinct $d$-dimensional volumes.
Department of Applied Mathematics,
Faculty of Mathematics and Physics, Charles University, Czech Republic
koki@kam.mff.cuni.cz
1 Introduction
For $d\geq 2$, a finite point set $X$ in $\mathbb{R}^{d}$ with $|X|\geq d+1$ is said to be in general position if no $d+1$ points of $X$ lie on the same hyperplane, and we say that an arrangement of hyperplanes $\mathcal{A}$ in $\mathbb{R}^{d}$ with $|\mathcal{A}|\geq d+1$ is in general position if every $d$ hyperplanes intersect at exactly one point and no $d+1$ hyperplanes have a common point.
Throughout the paper, we use $O_{d}$ to denote the standard big-$O$ notation in which the constant may depend on $d$. We use analogous conventions for $\Omega_{d}$ and $\Theta_{d}$. For $1\leq k\leq d$, we will refer to $k$-dimensional volume and $k$-dimensional simplex as $k$-volume and $k$-simplex. Furthermore, by “a $k$-simplex in a hyperplane arrangement in $\mathbb{R}^{d}$”, we mean a simplicial $k$-cell determined by $k+1$ hyperplanes of the arrangement.
1.1 Erdős’s distinct distances problem
In 1946, Erdős [erdos1946] posed the question: what is the minimum number $g_{d}(n)$ of distinct distances determined by $n$ points in $\mathbb{R}^{d}$. This question is known as Erdős’s distinct distances problem, and a variety of interesting results have been obtained. For the planar case $d=2$, Erdős showed in the same paper an upper bound $g_{2}(n)=O(n/\sqrt{\log n})$ by considering the $\sqrt{n}\times\sqrt{n}$ square lattice. While there had been earlier progress on lower bounds ([chung1984number, chung1992number, katz2004new, moser1952different, solymosi2001distinct, ekely1993crossing, tardos2003distinct]), a major breakthrough by Guth and Katz [guth2015erdHos] proved $g_{2}(n)=\Omega(n/\log n)$.
In higher dimensions, a $d$-dimensional grid $n^{1/d}\times\cdots\times n^{1/d}$ gives $g_{d}(n)=O_{d}(n^{2/d})$, and this bound is conjectured to be optimal. The best known lower bound to date is due to Solymosi and Vu [solymosi2008near] which is $g_{d}(n)=\Omega_{d}(n^{2/d-2/(d(d+2))})$.
A well-studied variant of this problem asks for the minimum number $g^{\prime}_{d}(n)$ of distinct $d$-volumes of $d$-simplices spanned by $n$ points in $\mathbb{R}^{d}$, assuming that not all $n$ points lie on a common hyperplane. One easily sees that $g^{\prime}_{d}(n)\leq\lfloor(n-1)/d\rfloor$ by taking $d$ sets of about $n/d$ equally spaced points on $d$ parallel lines passing through the vertices of a fixed (d1)(d-1)-simplex. Erdős, Purdy, and Straus [erdos1982problem] conjectured that this construction is tight for all sufficiently large $n$. Pinchasi [pinchasi2008minimum] proved this conjecture for the case $d=2$, and Dumitrescu and Cs.Tóth [dumitrescu2008number] later showed that $g^{\prime}_{d}(n)=\Theta_{d}(n)$ for every $d$.
There is another related problem asking (A): what is the maximum number $h_{d}(n)$ such that any set of $n$ points in $\mathbb{R}^{d}$ contains a subset $S$ of size $h_{d}(n)$ whose induced $d$-simplices all have distinct $d$-volumes. The upper bound is again derived from the $n^{1/d}\times\cdots\times n^{1/d}$ grid, which gives $h_{2}(n)=O_{d}(n^{2/3})$ and $h_{d}(n)=O_{d}(n^{(d-1)/d})$ for $d\geq 3$. Conlon et al. [conlon2015distinct] showed $h_{d}(n)=\Omega_{d}(n^{1/(2d+2)})$ for all $d\geq 2$.
1.2 Erdős’s unit distance problem
The unit distance problem asks: what is the maximum number $u_{d}(n)$ of unit distances that a set of $n$ points in $\mathbb{R}^{d}$ can form. For the case $d=2$, Erdős [erdos1946] obtained the bounds $n^{1+c/\log\log n}<u_{2}(n)=O(n^{3/2})$ for some $c>0$ and conjectured that his lower bound is tight. The problem remains wide open; the best known upper bound $u_{2}(n)=O(n^{4/3})$ is due to Spencer, Szemerédi, and Trotter [spencer1984unit]. Despite numerous efforts ([clarkson1990combinatorial, pach2006forbidden, ekely1993crossing]), no further improvements have been found. In $3$-dimensions, the best known bounds are $u_{3}(n)=O(n^{3/2})$ [kaplan2012unit, zahl2013improved] and $u_{3}(n)=\Omega(n^{4/3}\log\log n)$ [erdos1960set]. For $d\geq 4$, we have $u_{d}(n)=\Theta_{d}(n^{2})$ (see [erdHos1990variations, lenz1955zerlegung]).
This problem also has a well-studied direction of generalization. In 1967, Oppenheim asked (B): what is the maximum number of unit area triangles determined by $n$ points in the plane? See [erdHos1996extremal]. A simple lattice $\sqrt{\log n}\times(n/\sqrt{\log n})$ again gives a lower bound $\Omega(n^{2}\log\log n)$ [erdos1971some]. Currently, the best known upper bound is $O(n^{2+2/9})$, due to Raz and Sharir [raz2017number]. Dumitrescu and Cs.Tóth [dumitrescu2008number] extended this problem to three dimensions and obtained the following result: the number of unit volume tetrahedra ($3$-simplices) determined by $n$ points in $\mathbb{R}^{3}$ is $O(n^{7/2})$, and there is an $n$-point set achieving $\Omega(n^{3}\log\log n)$.
Related problems (C): are to determine the number of triangles of minimum/maximum area determined by a set of $n$ points studied by Braß, Rote, and Swanepoel [brass2001triangles]. They proved that the maximum number of triangles of maximum area determined by $n$ points in the plane is exactly $n$, and the maximum number of triangles of minimum area determined by $n$ points in the plane is $\Theta(n^{2})$.
Dumitrescu and Cs.Tóth [dumitrescu2008number] further generalized this problem and obtained the following results: The number of tetrahedra of minimum volume spanned by $n$ points in $\mathbb{R}^{3}$ is at most $2n^{3}/3-O(n^{2})$, and there is a configuration of $n$-points achieving $3n^{3}/16-O(n^{2})$. Furthermore, they proved that the maximum number of $d$-simplices of minimum volume spanned by $n$ points in $\mathbb{R}^{d}$ is $O_{d}(n^{d})$ for all $d$.
1.3 Dual setting
In this paper, we investigate the following dual setting of (A), (B) and (C).
  1. 1.
    The following is a dual setting of (A):
    Question 1.
    What is the maximum number $D_{d}(n)$ such that any arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$ in general position contains a subset of size $D_{d}(n)$ whose induced $d$-simplices have all distinct $d$-volumes?
    For the case $d=2$, Damásdi et al. [damasdi2020triangle] obtained a lower bound of $n^{1/5}$ under the stronger assumption that there are no six lines that are tangent to a common conic (note that five lines always have a common tangent conic).
  2. 2.
    The problem of generalizing (B) to higher dimensions is to determine the maximum number of unit $d$-volume $d$-simplices formed by $n$ points in $\mathbb{R}^{d}$. For this problem, we consider the following dual version.
    Question 2.
    What is the maximum number $f_{d}(n)$ of unit $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$?
    For the case $d=2$, Damásdi et al. [damasdi2020triangle] proved that $f_{2}(n)=O(n^{9/4+\varepsilon})$ for every fixed $\varepsilon>0$, and $f_{2}(n)=\Omega(n^{2})$. Their proof uses the incidence theorem between points and algebraic curves, proved by Sharir and Zahl [sharir2017cutting].
  3. 3.
    We propose the following variants for the higher-dimensional generalization of (C) (that is, to determinie the maximum number of $d$-simplices of minimum/maximum $d$-volume determined by a set of $n$ points in $\mathbb{R}^{d}$).
    Question 3.
    What is the maximum number $m_{d}(n)$ of minimum $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$ ?
    For the case $d=2$, Damásdi et al. [damasdi2020triangle] proved that $m_{2}(n)=\Theta(n^{2})$.
    Question 4.
    What is the maximum number $M_{d}(n)$ of maximum $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$ ?
    As we mentioned above, with respect to the occurrence of the maximum area determined by $n$ points in the plane, it is known that the maximum number is exactly $n$. However, Damásdi et al. [damasdi2020triangle] showed that this phenomenon does not occur in the dual setting: $M_{2}(n)>7n/5-O(1)$. We note that they also showed an upper bound $M_{2}(n)\leq n(n-1)/3$.
1.4 Our results
We provide partial answers to Questions 1-4, especially for $d\geq 3$.
Theorem 1.1.
For the maximum number $f_{d}(n)$ of unit $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
$f_{d}(n)=O_{d}\bigl(n^{\,d+1-\tfrac{d}{d^{d}+1}}\bigr).$
Theorem 1.2.
For the maximum number $f_{d}(n)$ of unit $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
$f_{d}(n)=\Omega_{d}(n^{d}).$
Theorem 1.3.
For the maximum number $m_{d}(n)$ of minimum $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
$m_{d}(n)=\Theta_{d}(n^{d}).$
Theorem 1.4.
For the maximum number $M_{d}(n)$ of maximum $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
$M_{d}(n)>\frac{7}{5}(n-d)-O(1).$
Theorem 1.5.
For the largest subset of hyperplanes forming $d$-simplices of distinct $d$-volumes, we have
$D_{2}(n)\ll n(\log n)^{-c}$
for some absolute constant $c>0$, For every $d\geq 3$, there is $c_{d}\in(0,1)$ such that
$D_{d}(n)\ll ne^{-(\log\log n)^{c_{d}}}.$
Theorem 1.5 gives partial answers for Problem 5.4 and Problem 5.5 in [damasdi2020triangle].
The rest of the paper is organized as follows. We will prove Theorem 1.1, 1.2 in the next section. In Section 3 we will prove Theorem 1.3. Section 4 is devoted to proving Theorem 1.4. Finally, in Section 5, we will give a proof of Theorem 1.5.
2 Unit volume $d$-simplices
In this section, we consider the number of $d$-simplices with unit volume determined by an arrangement of $n$ hyperplanes, namely to investigate the magnitude of $f_{d}(n)$. We have a trivial upper bound $O_{d}(n^{d+1})$ by selecting generic $d+1$ hyperplanes from the arrangement. The following provides a slightly better upper bound for $f_{d}(n)$. See 1.1
Proposition 2.1.
Let $H_{1},\dots,H_{d}$ be hyperplanes in $\mathbb{R}^{d}$ meeting at a unique point and fix $V_{0}>0$. Then, every hyperplane that form a $d$-simplex with $H_{1},\ldots,H_{d}$ of $d$-volume $V_{0}$ are all tangent to one of two fixed algebraic hypersurfaces of degree $d$ which have $H_{1},\ldots,H_{d}$ as asymptotes and each has exactly $2^{d-1}$ connected components. Moreover, the $2^{d}$ connected components of these two hypersurfaces are contained one each in the $2^{d}$ connected components determined by $H_{1},\ldots,H_{d}$.
Proof.
Note that an (invertible) affine transformation preserves hyperplanes, hypersurfaces of degree $d$ and ratios of volumes. Fix a coordinate $(x_{1},\ldots,x_{d})$ and we may assume $H_{i}=\{x_{i}=0\}\quad(i=1,\ldots,d).$ Fix $c\neq 0$ and consider the hypersurface
$S_{c}:=\Bigl\{x\in\mathbb{R}^{d}:\ \prod_{i=1}^{d}x_{i}=c\Bigr\}.$
Take an arbitrary point $p=(p_{1},\dots,p_{d})\in S_{c}$. Then $\prod_{i=1}^{d}p_{i}=c$, and $p_{i}\neq 0$ for all $i$. The tangent hyperplane of $S_{c}$ at $p$ is given by $\sum_{i=1}^{d}\frac{x_{i}}{p_{i}}=d.$ This tangent hyperplane together with $H_{1},\dots,H_{d}$ forms a $d$-simplex $\Delta_{p}$ with vertices
$O=(0,\dots,0),\quad(dp_{1},0,\dots,0),\quad(0,dp_{2},0,\dots,0),\ \dots,\ (0,\dots,0,dp_{d}).$
Hence the volume of $\Delta_{p}$ equals
$\frac{1}{d!}\prod_{i=1}^{d}|dp_{i}|=\frac{d^{d}}{d!}\,\Bigl|\prod_{i=1}^{d}p_{i}\Bigr|=\frac{d^{d}}{d!}\,|c|.$
Since the volume of $\Delta_{p}$ depends only on $d$ and $|c|$, for any point of $S_{-c}$, the tangent hyperplane at that point together with $H_{1},\ldots,H_{d}$ also cuts out a $d$-simplex of volume $\frac{d^{d}}{d!}\,|c|$. The hypersurfaces $S_{c}$ and $S_{-c}$ have $2^{d-1}$ connected components each, and every connected components has $H_{1},\ldots,H_{d}$ as asymptotes. Moreover, each of the $2^{d}$ regions determined by $H_{1},\ldots,H_{d}$ contains exactly one connected component of $S_{c}\cup S_{-c}$. Consequently, any hyperplane that, together with $H_{1},\ldots,H_{d}$, forms a $d$-simplex of volume $\frac{d^{d}}{d!}\,|c|$ is tangent to exactly one connected component of $S_{c}\cup S_{-c}$. One easily sees that the $2^{d}$ connected components of $S_{c}\cup S_{-c}$ are contained one each in the $2^{d}$ connected components determined by $H_{1},\ldots,H_{d}$. ∎
Proof of Theorem 1.1.
The proof is based on [dumitrescu2009extremal]. Let $\mathcal{H}$ be an arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$. For every $d$-tuple $(H_{i_{1}},\ldots,H_{i_{d}})$ of hyperplanes in $\mathcal{H}$ with $\bigcap_{j=1}^{d}H_{i_{j}}$ consisting of a single point, apply the above Proposition with $V_{0}=1$ to obtain the two algebraic hypersurfaces of degree $d$. Let $\sum\nolimits_{d}$ be the union of all such connected components by choosing $d$ hyperplanes intersecting at a single point. Since each hypersurface has $2^{d-1}$ connected components, we have $|\sum\nolimits_{d}|\leq 2\cdot 2^{d-1}\binom{n}{d}$.
We consider the following dual transformation in $\mathbb{R}^{d}$ between hyperplanes $H:x_{d}=p_{1}x_{1}+p_{2}x_{2}+\cdots+p_{d-1}x_{d-1}+p_{d}$ and points $H^{*}=(p_{1},p_{2},\ldots,p_{d-1},p_{d})$. For a connected component $\sigma\in\sum\nolimits_{d}$, define $X_{\sigma}:=\{H^{*}:H\text{ is tangent to }\sigma\}.$ Then, $X_{\sigma}$ is contained in an algebraic hypersurface of degree at most $d$. Moreover, $X_{\sigma}$ has at most two connected components: it has two connected components if and only if $\sigma$ admits a tangent hyperplane parallel to the $x_{d}$-axis.
Now we consider two bipartite graphs $G_{1},G_{2}$ with vertex set $\mathcal{H}\sqcup\sum\nolimits_{d}$ as follows. If $H\in\mathcal{H}$ is tangent to $\sigma\in\sum\nolimits_{d}$ from above (in the sense of the last axis), we define an edge $(H,\sigma)\in\mathcal{H}\times\sum\nolimits_{d}$ in $G_{1}$. Similarly, If $H\in\mathcal{H}$ is tangent to $S\in\sum\nolimits_{d}$ from below, we define an edge $(H,\sigma)\in\mathcal{H}\times\sum\nolimits_{d}$ in $G_{2}$. The edges of $G_{1}$ and $G_{2}$ correspond to all hyperplane-hypersurface tangencies. Hence, to obtain an upper bound for $f_{d}(n)$, it suffices to bound the number of edges $E(G_{1})$ in $G_{1}$ and the number of edges $E(G_{2})$ in $G_{2}$. Observe that neither $G_{1}$ nor $G_{2}$ contains $K_{d^{d}+1,d}$ as a subgraph. Indeed, if $G_{1}$ contains $K_{d^{d}+1,d}$, then there would be $d^{d}+1$ hyperplanes in $\mathcal{H}$ tangent to $d$ connected components from above. Considering this situation in a dual space, meaning that $d$ connected components intersect at $d^{d}+1$ points, which contradicts the higher-dimensional Bézout’s theorem (refer to [tao2011bezout]). The same argument applies to $G_{2}$. Therefore, by the Kövári–Sós–Turán theorem,
$|E(G_{1})|,|E(G_{2})|=O_{d}\bigl(n\cdot n^{d(1-\frac{1}{d^{d}+1})}+n^{d}\bigr)=O_{d}\bigl(n^{\,d+1-\tfrac{d}{d^{d}+1}}\bigr).$
It is clear that a lower bound for $m_{d}(n)$ gives a lower bound for $f_{d}(n)$. See the lower bound construction in the next section.
3 Minimum volume $d$-simplices
In this section, we prove Theorem 1.3. Let us recall the statement. See 1.3
Kobon Fujimura asked in his book (called ”The Tokyo puzzle” [Fujimura1978]) what is the largest number $K(n)$ of non-overlapping triangles whose sides lie on an arrangement of $n$ lines (Grünbaum [grunbaum1972arrangements] had studied a similar problem slightly earlier than Fujimura, but only in the case of arrangements in the projective plane). A first upper bound was established by Saburo Tamura, which is $K(n)\leq\lfloor\frac{n(n-2)}{3}\rfloor$. In 2007, Clément and Bader [clement2007tighter] found a slightly improved upper bound:
$K(n)\leq\begin{cases*}\lfloor\frac{n(n-2)}{3}\rfloor-1&if $n\equiv 0,2\pmod{6}$,\\ \lfloor\frac{n(n-2)}{3}\rfloor&otherwise.\end{cases*}$
For even $n$, Bartholdi et al. [bartholdi2008simple] provided a slightly sharper upper bound of $\lfloor\frac{n}{3}(n-\frac{7}{3})\rfloor$. On the other hand, Füredi and Palásti [furedi1984arrangements] having constructed an arrangement of $n$ lines which contains $\lfloor\frac{n(n-3)}{3}\rfloor$ non-overlapping triangles, which gives $K(n)\geq\lfloor\frac{n(n-3)}{3}\rfloor$.
We now consider a higher-dimensional analogue. Let $K_{d}(n)$ be the largest number of nonoverlapping $d$-simplices whose facets lie on an arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$. Note that $K(n)=K_{2}(n)$.
For $d=2$, it immediately follows that $m_{2}(n)\leq K_{2}(n)$ from the fact that cutting a triangle with a single line always forms a smaller triangle in its interior. However, unfortunately, the same argument does not hold for $d\geq 3$. Indeed, a $d$-simplex cut by a hyperplane creates a new $d$-simplex with $d$-volume smaller than the original in the interior if and only if the hyperplane separates one vertex from all the others. See Figure 1 for the case $d=3$.
Figure 1: Left: a cut creating a $3$-simplex in the interior of a 33-simplex. Right: a cut not creating a $3$-simplex in the interior.
Nevertheless, finding bounds for $K_{d}(n)$ might be an interesting problem for its own sake.
Proposition 3.1.
For any $d\geq 3$, and $n\geq d+1$,
$K_{d}(n)\leq\frac{2}{(d+1)!}n^{d}+O_{d}(n^{d-1}).$
The proof is presented in the Appendix. We provide an upper bound for $m_{d}(n)$ by a different approach.
Proof of Theorem 1.3.
Upper bound: Let $\mathcal{H}=\{H_{1},\dots,H_{n}\}$ be an arrangement of hyperplanes in $\mathbb{R}^{d}$, and let $m_{0}$ be the minimum $d$-volume of a $d$-simplex determined by $\mathcal{H}$. Assume that there exists a $d$-tuple $(H_{i_{1}},\dots,H_{i_{d}})$ of hyperplanes in $\mathcal{H}$ meeting at exactly one point. Applying Proposition 2.1 to $(H_{i_{1}},\dots,H_{i_{d}})$ with $V_{0}=m_{0}$, any hyperplane $T$ such that $\{H_{i_{1}},\dots,H_{i_{d}},T\}$ forms a $d$-simplex of $d$-volume $m_{0}$ must be tangent to one of the two fixed algebraic hypersurfaces of degree $d$. These two hypersurfaces have $2^{d}$ connected components in total, and each of the $2^{d}$ connected components determined by $H_{i_{1}},\dots,H_{i_{d}}$ contains exactly one of them.
Let $C_{R}$ be the unique connected component contained in a fixed connected component $R$ in $\{H_{i_{1}},\dots,H_{i_{d}}\}$ . It suffices to show that at most one hyperplane in $\mathcal{H}\setminus\{H_{i_{1}},\dots,H_{i_{d}}\}$ is tangent to $C_{R}$. For a contradiction, we assume that there exist two distinct hyperplanes $T,T^{\prime}\subset\mathcal{H}\setminus\{H_{i_{1}},\dots,H_{i_{d}}\}$ tangent to $C_{R}$. Let $\tau$ be the $d$-simplex formed by $H_{i_{1}},\dots,H_{i_{d}},T$, and let $\tau^{\prime}$ be the $d$-simplex formed by $H_{i_{1}},\dots,H_{i_{d}},T^{\prime}$. Note that $\tau$ and $\tau^{\prime}$ are $d$-simplices of $d$-volume $m_{0}$.
Now consider the arrangement $\mathcal{A}=\{H_{i_{1}},\dots,H_{i_{d}},T,T^{\prime}\}$. A well-known result of Shannon [shannon1979simplicial] states that an arrangement of $N(\geq d+1)$ hyperplanes in general position in $\mathbb{R}^{d}$ has NdN-d simplicial $d$-cells, and hence $\mathcal{A}$ has two simplicial $d$-cells. Observe that these two simplicial $d$-cells have $T$ as a facet, and they do not lie on the same side of $T$. Therefore, at least one of them is strictly contained in $\tau$ or $\tau^{\prime}$. This contradicts the minimality of $m_{0}$. Hence, we have $m_{d}(n)\leq 2^{d}\binom{n}{d}=O_{d}(n^{d})$.
Lower bound: Denote by $S_{d}$ the set of all permutations of $\{1,\ldots,d\}$ and let $\pi\in S_{d}$ be a permutation. The Coxeter–Freudenthal–Kuhn (CFK) triangulation of the unit hypercube $0,1]^{d}$ is a triangulation of $0,1]^{d}$ into $d$-simplices $\sigma_{\pi}=[v_{0}(\pi),\ldots,v_{d}(\pi)]$, whose vertices are given by
$v_{0}(\pi)=(0,\ldots,0),\quad v_{i}(\pi)=v_{i-1}(\pi)+e_{\pi(i)}\quad\text{for }i=1,\ldots,d,$
where $e_{1},\ldots,e_{d}$ is a standard basis of $\mathbb{R}^{d}$. Note that the points $v_{0}(\pi),\ldots,v_{d}(\pi)\in\mathbb{R}^{d}$ are vertices of $0,1]^{d}$ and the CFK triangulation decomposes $0,1]^{d}$ into $d!$ $d$-simplices with $d$-volume $1/d!$. It is easy to observe that this decomposition is realized by $\binom{d}{2}$ hyperplanes $\{x_{p}-x_{q}=0:1\leq p<q\leq d\}.$ See Figure 2 and refer to [coxeter1934discrete], [freudenthal1942simplizialzerlegungen], [kuhn1968simplicial] for further details
(1,1,1)(1,1,1)(0,0,0)(0,0,0)(0,0)(0,0)$(1,1)$
Figure 2: The CFK triangulation of $0,1]^{2}$ and $0,1]^{3}$.
Now, we consider $d(n+1)$ hyperplanes given by $x_{i}=k\quad(i\in\{1,\ldots,d\},\,\,k\in\{0,\ldots,n\}),$ which form an arrangement that decomposes the hypercube $0,n]^{d}$ into $n^{d}$ unit hypercubes. Furthermore, by adding $\binom{d}{2}\cdot(2n-1)$ hyperplanes
$\{x_{p}-x_{q}=t:1\leq p<q\leq d,\,t\in\mathbb{Z},\,|t|\leq n-1\},$
We can see that each unit hypercube admits the CFK triangulation (this is called K1 triangulation [REIA2025104237]). The size of the resulting hyperplane arrangement is
$d(n+1)+\binom{d}{2}(2n-1)=d^{2}n+\frac{d(3-d)}{2},$
and the number of (minimum volume) $d$-simplices is $d!\cdot n^{d},$ which gives us a lower bound $\Omega_{d}(n^{d})$.
4 Maximum volume tetrahedra
In this section, we show a lower bound for $M_{d}(n)$. A linear lower bound $M_{d}(n)\geq n-d$ can be easily obtained as follows: Fix coordinates $(x_{1},\dots,x_{d})$ in $\mathbb{R}^{d}$ and let $H_{i}:=\{x_{i}=0\}$, $i=1,\dots,d$ and consider the branch
$\sigma:=\{x\in\mathbb{R}^{d}:\ \prod_{i=1}^{d}x_{i}=1,\ x_{1},\ldots,x_{d}>0\}.$
Choose ndn-d distinct hyperplanes $T_{1},\dots,T_{n-d}$ tangent to $\sigma$, with their tangent points contained in a sufficiently small neighborhood of some fixed point $u\in\sigma$. For each $j$, the hyperplanes $H_{1},\dots,H_{d},T_{j}$ bound a $d$-simplex in the orthant $\{(x_{1},\ldots,x_{d}):x_{1},\ldots,x_{d}\geq 0\}$, and its $d$-volume equals $d^{d}/d!$. Since the points of tangency of $T_{1},\dots,T_{n-d}$ are chosen sufficiently close to each other, any bounded $d$-simplex of the arrangement that uses at least two of the hyperplanes $T_{j}$ has $d$-volume smaller than $d^{d}/d!$. Hence the simplices bounded by $H_{1},\dots,H_{d}$ and one $T_{j}$ are of maximum $d$-volume in the arrangement, and therefore $M_{d}(n)\geq n-d$.
4.1 Explicit lower bound construction for $M_{3}(n)$
As mentioned in the introduction, Damasdi et al. [damasdi2020triangle] obtained the lower bound $M_{2}(n)>\frac{6}{5}n-O(1)$. The idea of their lower bound construction is the following: They begin with a family of five lines forming a regular pentagon, and then recursively place arrangements together. At each step, two arrangements are combined by applying a suitable affine transformation so that the areas of their maximum area triangles coincide, and placing them in such a way that the maximum area does not increase. Moreover, by sliding one of the whole families of lines, one can create two additional maximum area triangles. Consequently, we obtain a construction in which the number of maximum area triangles increases by seven whenever a family of five lines is added. We see that this idea can also work in three dimensions.
Proposition 4.1.
$M_{3}(n)>\frac{7}{6}n-O(1).$
The following can be proved using the same procedure as in their proof. Details are left to the reader.
Proposition 4.2.
If there are no two parallel planes in the arrangement, then we can find a rectangular box $ABCDEFGH$ such that adding any plane to the arrangement that intersects the segments $AB$, CDCD, EFEF and $GH$ creates no new maximum volume tetrahedra.
For a plane arrangement $\mathcal{A}$, let $T(\mathcal{A})$ denote the number of maximum volume tetrahedra in $\mathcal{A}$.
$A_{1}$$B_{1}$$D_{1}$$C_{1}$$H_{1}$$G_{1}$$E_{1}$$F_{1}$$A_{2}$$B_{2}$$C_{2}$$D_{2}$$E_{2}$$F_{2}$$G_{2}$$H_{2}$
Figure 3: Rectangular boxes
Figure 4: A star-shaped badge
Proposition 4.3 ($3$-dimensional analogue of Proposition 4.2 in [damasdi2020triangle]).
If $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are arrangements of planes neither of which contains two parallel planes, then there exist affine transformations $\varphi,\psi$ such that
$T(\varphi(\mathcal{H}_{1})\cup\psi(\mathcal{H}_{2}))\geq T(\mathcal{H}_{1})+T(\mathcal{H}_{2})+2.$
Proof.
In an arrangement $\mathcal{H}_{1}$, there exists a rectangular box satisfying Proposition 4.2 with vertices $A_{1},B_{1},C_{1},D_{1},E_{1},F_{1},G_{1},H_{1}$, and similarly in $\mathcal{H}_{2}$, there exists one with vertices $A_{2},B_{2},C_{2},D_{2},E_{2},F_{2},G_{2},H_{2}$. By applying appropriate affine transformations to each, we can arrange $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ so that the volumes of their respective maximum volume tetrahedra are equal and the two rectangular boxes cross each other as shown in Figure 4. Let $\varphi$ denote the affine transformation applied to $\mathcal{H}_{1}$ and $\psi$ the affine transformation applied to $\mathcal{H}_{2}$.
As the next step, we translate the planes of $\psi(\mathcal{H}_{2})$ to create a new maximum-volume tetrahedron $S$. If exactly one of the four planes defining $S$ comes from a plane of $\varphi(\mathcal{H}_{1})$, let this plane be $H$. By moving the entire $\psi(\mathcal{H}_{2})$ along $H$, we can create another new maximum-volume tetrahedron. Similarly, if exactly three planes come from planes of $\varphi(\mathcal{H}_{1})$, we can move $\varphi(\mathcal{H}_{1})$ along one plane of $\psi(\mathcal{H}_{2})$. Finally, if exactly two planes come from planes of $\varphi(\mathcal{H}_{1})$, by moving $\psi(\mathcal{H}_{2})$ along the line of intersection of those two planes, we can create a new maximum-volume tetrahedron. ∎
We observe that the arrangement of planes $\mathcal{S}\subset\mathbb{R}^{3}$ forming a star-shaped badge as shown in Figure 4 consists of six planes and forming five maximum volume tetrahedra. Starting from $\mathcal{S}$, we repeatedly add suitably transformed copies of $\mathcal{S}$ so that the above proposition applies at each step. Consequently, for every $k\geq 1$ we obtain an arrangement of $6k$ planes in $\mathbb{R}^{3}$ that determines at least 7k27k-2 maximum volume tetrahedra, which proves Theorem 1.4.
4.2 General lower bound for $M_{d}(n)$
We improve here the lower bound achieved by Proposition 4.1 and also provides lower bounds for all $d\geq 3$.
See 1.4
The Theorem follows immediately from the following observation.
Observation 1.
For all $d\geq 3$, we have
$M_{d}(n)\geq M_{d-1}(n-1).$
Proof.
[…]
We are not aware of any upper bound for $M_{d}(n)$ that holds for all $d\geq 3$. Moreover, we do not have any reason to believe that the correct upper bound is also close to linear. Whether $M_{d}(n)$ is linear in $n$ is open even in the case $d=2$. Note that Damásdi et al. [damasdi2020triangle] conjectured that $M_{2}(n)=O(n^{1+\varepsilon})$ for every $\varepsilon>0$.
5 Hyperplanes defining distinct volume $d$-simplices
In this section, we show that for all $d\geq 2$, $D_{d}(n)$ grows strictly slower than $n$. See 1.5
Let $n]=\{1,\ldots,n\}$ and denote by $r_{k}(n)$ the size of the largest $S\subseteq[n]$ such that $S$ contains no $k$ elements in arithmetic progression. The following is the key proposition.
Proposition 5.1.
$D_{d}(n)<r_{d+2}(n)$ for all $d\geq 2$.
Determining the value of $r_{k}(n)$ is one of the fundamental problems in arithmetic combinatorics, and both upper and lower bounds have been progressively improved. In 2009, Green and Tao [green2006new, green2017new] proved that there exists a constant $c>0$ such that $r_{4}(n)\ll n(\log n)^{-c}$. For $k\geq 4$, Leng et al. [leng2024improved] recently obtained an upper bound $r_{k}(n)<<ne^{-(\log\log n)^{c_{k}}}$ for some $c_{k}\in(0,1)$. See [szemeredi1975sets, gowers1998new, gowers2000arithmetic] for further details. Theorem 1.5 can be immediately derived from these bounds and Proposition 5.1. Therefore, any improvement of upper bound for $r_{k}(n)$ automatically improves that for $D_{d}(n)$.
Proof of Proposition 5.1.
We first consider the case $d=2$. Define $\mathcal{L}=\{l_{1},\dots,l_{n}\}$ as the set of lines corresponding to $n$ consecutive edges of a regular $N$-gon $(N>n)$, arranged in clockwise order. We choose $N$ sufficiently large so that any four lines in $\mathcal{L}$ form two triangles. By the definition of $r_{k}(n)$, every subset of $\mathcal{L}$ of size $r_{4}(n)+1$ always contains four lines $\{l_{i_{1}},l_{i_{2}},l_{i_{3}},l_{i_{4}}\}$ such that $\{i_{1},i_{2},i_{3},i_{4}\}\subseteq[n]$ is a $4$-terms arithmetic progression. One can easily see that these four lines then form two congruent triangles.
We generalize this construction to higher dimensions. First, assume that $d$ is even. Write $\mathbb{R}^{d}$ as the orthogonal sum of two-dimensional subspaces:
$\mathbb{R}^{d}=V_{1}\oplus\cdots\oplus V_{d/2}.$
For each $j$, let $C_{j}$ denote the unit circle centered at the origin in $V_{j}$. We choose angle parameters $\theta_{1},\ldots,\theta_{d/2}>0$ so that for each $j$, one has $n\theta_{j}\ll\pi/2,\,\text{and}\,\,\theta_{j}\neq\theta_{j^{\prime}}\ \text{for }j\neq j^{\prime}.$ For each $i\in\{0,\ldots,n\}$ and each $j\in\{1,\ldots,d/2\}$, we define $P_{i,j}=(\cos(i\theta_{j}),\sin(i\theta_{j}))\in V_{j}$ (viewed in the standard coordinates of $\mathbb{R}^{d}$, the point $P_{i,j}$ has nonzero coordinates only in its $V_{j}$-component, which is $(\cos(i\theta_{j}),\sin(i\theta_{j}))$, while all other components are zero). For each $i\in\{1,\ldots,n\}$, let $H_{i}$ denote the affine hull of
$\bigcup_{1\leq j\leq d/2}\{P_{i-1,j},P_{i,j}\}.$
This has dimension d1d-1. Let $\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ denote the resulting hyperplane arrangement. One can verify that $\mathcal{H}$ is in general position for almost all choices of the angle parameters $\theta_{1},\ldots,\theta_{d/2}$. For any positive integer $D$, we define the map $T_{D}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ as
$T_{D}|_{V_{j}}=R_{D\theta_{j}}\,(j=1,\ldots,d/2),$
where $R_{D\theta_{j}}$ is the rotation by angle $D\theta_{j}$ on $V_{j}$. One easily sees that $T_{D}(P_{i,j})=P_{i+D,j}$ (for $i+D\leq n$), and therefore $T_{D}(H_{i})=H_{i+D}.$ Since the Jacobian matrix of $T_{D}$ is
$J_{T_{D}}=\begin{pmatrix}R_{D\theta_{1}}&0&\cdots&0\\ 0&R_{D\theta_{2}}&\ddots&\vdots\\[2.0pt] \vdots&\ddots&\ddots&0\\[2.0pt] 0&\cdots&0&R_{D\theta_{d/2}}\end{pmatrix}$,
$\det J_{T_{D}}=\prod_{j=1}^{d/2}\det R_{D\theta_{j}}=1.$ Thus, the map $T_{d}$ is an affine volume-preserving map.
Now let $i_{1}<i_{2}<\cdots<i_{d+2}$ be a $(d+2)$-arithmetic progression in $[n]$ with common difference $L$. As the arrangement $\mathcal{H}$ is in general position, the hyperplanes $\{H_{i_{1}},\ldots,H_{i_{d+1}}\}$ form a unique $d$-simplex $S$, and the hyperplanes $\{H_{i_{2}},\ldots,H_{i_{d+2}}\}$ form a unique $d$-simplex $S^{\prime}$. Since $T_{L}(H_{i_{k}})=H_{i_{k+1}}$, we have $T_{L}(S)=S^{\prime}$ which means that $S$ and $S^{\prime}$ have the same $d$-volume. Therefore, by the definition of $r_{k}(n)$, every subfamily of $\mathcal{H}$ of size (at least) $r_{d+2}(n)+1$ always contain two $d$-simplices with the same $d$-volume.
Now, we assume $d$ is odd, we decompose $\mathbb{R}^{d}$ into an orthogonal direct sum of (d1)/2(d-1)/2 2-dimensional vector spaces and one 1-dimensional vector space:
$\mathbb{R}^{d}=V_{1}\oplus\cdots\oplus V_{(d-1)/2}\oplus L.$
First, we place $n$ equally spaced points $Q_{i}=(0,\ldots,0,i)$ ($i\in\{1,\ldots,n\}$) on $L$. Choose angular parameters $\theta_{1},\ldots,\theta_{(d-1)/2}>0$ such that $\theta_{j}\neq\theta_{j^{\prime}}$ for $j\neq j^{\prime}$. For each $j\in\{1,\ldots,(d-1)/2\}$, consider the helical curve wrapping around $L$ in the $3$-dimensional vector space $V_{j}\oplus L$:
$\gamma_{\theta_{j}}(t)=(0,\ldots,0,\underbrace{\cos(t\theta_{j})}_{2j-1},\underbrace{\sin(t\theta_{j})}_{2j},0,\ldots,0,t)$
and we place $(n+1)$ equally spaced points on $\gamma_{\theta_{j}}(t)$:
$P_{i,j}=(0,\ldots,0,\underbrace{\cos(i\theta_{j})}_{2j-1},\underbrace{\sin(i\theta_{j})}_{2j},0,\ldots 0,i),\quad i\in\{0,\ldots,n\}.$
Define $H_{i}$ as the affine hull of
$\bigcup_{1\leq j\leq(d-1)/2}\{P_{i-1,j},P_{i,j}\}\cup\{Q_{i}\},$
which has dimension d1d-1. By taking $\theta_{1},\ldots,\theta_{(d-1)/2}$ appropriately, we obtain the hyperplane arrangement $\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ which is in general position. For any positive integer $D$, one can define an affine volume-preserving map $T_{D}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ by
$T_{D}|_{V_{j}}=R_{D\theta_{j}},\,T_{D}|_{L}:m\mapsto m+D.$
Observe that $T_{D}(P_{i,j})=P_{i+D,j},\,T_{D}((0,\ldots,0,i))=(0,\ldots,0,i+D)$ (for $i+D\leq n$), hence $T_{D}(H_{i})=H_{i+D}.$ The same argument as in the previous case shows that any subfamily of $\mathcal{H}$ with size $r_{d+2}(n)+1$ contains two $d$-simplices with the same $d$-volume.
Remark 1.
As previouly mentioned, Damásdi et al. [damasdi2020triangle] obtained the lower bound with some additional assumptions: if there are no six lines that are tangent to a common conic (observe that five lines always have a common tangent conic), then an arrangement of $n$ lines always contains a subset of size $\Omega(n^{1/5})$ whose induced triangles have all distinct $d$-volumes. Their argument can be applied to general dimensions, and the following can be shown: If there are no $\binom{2d}{d}$ hyperplanes that are tangent to a common hypersurface, then then an arrangement of $n$ hyperplanes always contains a subset of size $\Omega_{d}(n^{\frac{1}{2d+1}})$ whose induced $d$-simplices have all distinct volumes. However, We are not aware of any lower bounds for $D_{d}(n)$ for $d\geq 2$.
Acknowledgement. I would like to thank Pavel Valtr for helpful suggestions. The work was supported by grant no. 23-04949X of the Czech Science Foundation (GAČR).
Appendix
Proof of Proposition 3.1.
It suffices to show that
$K_{d}(n)\leq\frac{n}{d+1}K_{d-1}(n-1)$
holds for all $d\geq 3$. Suppose there are two parallel hyperplanes in an arrangement. If we perturb one of them by slightly rotating its normal vector, the number of simplicial $d$-cells in the arrangement does not decrease. Hence, we may assume that no two hyperplanes are parallel.
Consider an arrangement $\mathcal{H}$ of $n$ hyperplanes, and fix one hyperplane $H$ of $\mathcal{H}$. Then, restricting the remaining to $H$ yields an arrangement $\mathcal{S}$ of n1n-1 (d2)(d-2)-flats in $H(\cong\mathbb{R}^{d-1})$. Since no two of hyperplanes in $\mathcal{H}$ are parallel, no two in $\mathcal{S}$ are parallel either. Since the number of simplicial (d1)(d-1)-cells in $\mathcal{S}$ is at most $K_{d-1}(n-1)$, the total number of simplicial (d1)(d-1)-cells in $\mathcal{H}$ is at most $nK_{d-1}(n-1)$.
We now divide into two cases. First assume that no $d+1$ hyperplanes intersect at a single point. Then, no two simplicial $d$-cells share a common (d1)(d-1)-face. Otherwise, at least one of the $d$ vertices of the common face intersects at least $d+1$ hyperplanes. See Figure 6. Therefore,
$(d+1)K_{d}(n)\leq nK_{d-1}(n-1).$
Next, assume that $d+1$ (or more) hyperplanes intersect at a single point. If a $(d+1)$-th hyperplane is added through an existing intersection point, the number of simplicial (d1)(d-1)-faces decreases by $d+1$. On the other hand, two simplicial $d$-cells sharing a common simplicial (d1)(d-1)-face can be created only by adding a new hyperplane passing through an existing point. See Figure 6. Therefore,
$(d+1)K_{d}(n)<nK_{d-1}(n-1).$
Figure 5: The case $d=3$: if two tetrahedron share a common face, then at least one vertex of that face is contained in at least 44 planes.
Figure 6: The case $d=3$: if four planes intersect at a single point, adding a fourth plane through an existing point decreases the number of triangles by four, with at most two of them are saved.